OpenStudy (anonymous):

1/64=16^4x+3 solve for x what about this?

OpenStudy (angela210793):

2^(-6)=2^[4*(4x+3)] 16x+12=-6 16x=-18 x=-18/16 x=-9/8

OpenStudy (anonymous):

i cant get wht u are doing in the first step...

OpenStudy (angela210793):

1/64=1/(2^6)=2^(-6) 16=2^4 so 16^(4x+3)=2^[4(4x+3)]

OpenStudy (angela210793):

i have expressed both parts of the eq with base 2 so that i can equal exponents

OpenStudy (angela210793):

(x^a)^b=X^(ab) and 1/x^a=X^(-a)

OpenStudy (dumbcow):

If the two sides have same base, set the exponents equal to each other

OpenStudy (anonymous):

hold on let me try another and get back to u to check it

OpenStudy (angela210793):

kk